OpenStudy (kiiraa_x3):

When the polynomial P(x) = 6x3 + kx2 − x + 6 is divided by x − 1, the remainder is 0. Which of the following is also a factor of P(x)? 3x − 2 2x + 3 x + 1 3x + 2 HELPPPPPPPPP

OpenStudy (kiiraa_x3):

Solve 6x3 − x2 − 27x − 20 = 0 by factoring.

OpenStudy (aravindg):

According to remainder theorem x-1 is a factor Find the other factor.

OpenStudy (kiiraa_x3):

idk that's why I'm asking

jimthompson5910 (jim_thompson5910):

if x-1 is a factor of P(x), then P(1) = 0 P(x) = 6x^3 + kx^2 - x + 6 P(1) = 6(1)^3 + k(1)^2 - (1) + 6 ... replace every x with 1 0 = 6(1)^3 + k(1)^2 - (1) + 6 ... replace P(1) with 0 now solve for k